OpenStudy (anonymous):

Solve the problem as directed. The volume (v) of a sphere varies directly as the cube of its diameter (d). Write this statement in algebraic language, using an equation with the variables c, v, and d. C=???

OpenStudy (anonymous):

The volume (v) of a sphere varies directly as the cube of its diameter (d) Direct variation means v= c d^3 where c is a constant. and d is the diameter If you know the formula for the volume of a sphere (see above post) V = (4/3) * pi * r^3 or (replace r with d/2) V= ( pi/ 6 ) * d ^3 match V= ( pi/ 6 ) * d ^3 with ^ | V = c * d^3 to find the exact value for c

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

V = (4/3) * pi * r^3 V = (4/3) * pi * (d/2) ^ 3 V = (4/3) * pi * d^3/ 8 V =( pi/ 6 ) * d ^3

OpenStudy (superdavesuper):

C should be a constant; not a variable in the equation. You can find the value of C by using the formula: Volume of Sphere = (4/3)*pi*(Radius)^3 where Radius = Diameter/2.

OpenStudy (anonymous):

I don't understand